OpenStudy (anonymous):

Please Help!!!! I am really stuck on this one I used the formula m=y2-y1/x2-x1 and got -2/-3 but I am not sure how to put it into slope intercept form like the answers.

OpenStudy (anonymous):

https://www.desmos.com/calculator :D

OpenStudy (texaschic101):

Can you post the entire question

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (2,7)...x1 = 2 and y1 = 7 (-1,5)..x2 = -1 and y2 = 5 now sub slope(m) = (5 - 7) / (-1 - 2) slope(m) = -2/-3 slope(m) = 2/3 Now use y = mx + b form...we already know m, so now we need to find b, the y intercept. slope(m) = 2/3 you can use either of your points... (2,7)...x = 2 and y = 7 now sub and solve for b, the y intercept 7 = 2/3(2) + b 7 = 4/3 + b 7 - 4/3 = b 21/3 - 4/3 = b 17/3 = b -- turn to mixed number 5 2/3 = b so our equation is : y = 2/3x + 5 2/3 (answer B)

OpenStudy (texaschic101):

any questions...please ask :)

OpenStudy (anonymous):

why did the slope go from -2/-3 to regular 2/3?

OpenStudy (texaschic101):

because a negative divided by a negative is a positive

OpenStudy (anonymous):

I didn't know you were supposed to divide them, I thought it just stayed a fraction. Thank you for your help!! I fanned you :D

OpenStudy (texaschic101):

glad I could help :)